# Carleman estimates for third order operators of KdV and non KdV-type and applications

### 03 Jan 2024 v1

@@ 2. Carleman estimate and local solvability for $P_{1}$ and $P_{1}^{*}$
    $[X_{j},X_{k}](x,D)=\sum_{l=1}^{N}c^{jk}_{l}(x)X_{l}(x,D),\quad\forall x\in\mathbb{R}^{n+1}.$ (2.4)
@@
Notations. Below $N$ will always be a positive integer, $\lambda$ will be a real number, whereas $A=\{a_{ji}(x)\}_{\begin{subarray}{c}j=0,\ldots,N\\ i=1,\dots,n+1\end{subarray}}$ will denote the variable coefficient matrix containing the coefficients of the vector fields $iX_{j}$, for $j=0,\ldots,N$. For a matrix $A$ as above we define
$\|\nabla^{\alpha}A\|_{L^{\infty}(\Omega)}:=\sup_{\begin{subarray}{c}j=1,\ldots,N\\ i=1,\dots,n+1\end{subarray}}\|\partial^{\alpha}a_{ji}\|_{L^{\infty}(\Omega)},\quad\forall\Omega\subset\mathbb{R}^{n+1}.$ (2.8)
Throughout the paper we will write $\|\cdot\|_{L^{\infty}}$, $\|\cdot\|_{L^{2}}$, $\|\cdot\|_{H^{s}}$ without specifying the set $\Omega$ where the $L^{\infty}(\Omega)$, $L^{2}(\Omega)$, and the $H^{s}(\Omega)$ norms are taken. Given that we will be working with compactly supported functions, the set $\Omega$ will be the fixed support of the function (or its interior), so we omit it for simplicity. However, note that we can take $\|\nabla^{\alpha}A\|_{L^{\infty}}=\|\nabla^{\alpha}A\|_{L^{\infty}(\mathbb{R}^{n+1})}$ in the coming estimates when $A$ has $C^{\infty}_{b}(\mathbb{R}^{n+1})$ coefficients. Finally, we shall use the notation $\Gamma(\{X_{j}\}_{j=1}^{N})$ for the distribution spanned by the system of vector fields $\{X_{j}\}_{j=1}^{N}$.
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)$ (2.9)
$\geq\small{\sum_{j=1}^{N}}(Q_{j}e^{\lambda f}u,e^{\lambda f}u)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}e^{\lambda f}u,e^{\lambda f}u)-\lambda^{2}C_{K}\|e^{\lambda f}u\|_{L^{2}}^{2},$ (2.10)
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u).$ (2.13)
@@
$e^{\lambda f}P_{1}(x,D)u=e^{\lambda f}P_{1}(x,D)e^{-\lambda f}e^{\lambda f}u=P^{f}_{1}(x,D)e^{\lambda f}u,$ (2.14)
@@
where $L_{2}.L_{1},L_{0}$ are the partial differential differential operators of order 0, 1, and 2, given, respectively, by
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)$ $=\mathrm{Im}(P_{1}e^{\lambda f}u,e^{\lambda f}u)+\mathrm{Im}(L_{2}e^{\lambda f}u,e^{\lambda f}u)+\mathrm{Im}(L_{1}e^{\lambda f}u,e^{\lambda f}u)+\mathrm{Im}(L_{0}e^{\lambda f}u,e^{\lambda f}u),$ (2.29)
so, for simplicity, we handle each term on the left hand side separately.
@@
Now, on putting $v:=e^{\lambda f}u$, and using the previous observation, we have
$2i\mathrm{Im}(P_{1}e^{\lambda f}u,e^{f}u)$ $=2i\sum_{j=1}^{N}\mathrm{Im}(X_{1}X_{j}^{*}X_{j}v,v)+2i\,\mathrm{Im}(X_{0}v,v)$ (2.32)
$=\sum_{j=1}^{N}((X_{1}X_{j}^{*}X_{j}-X_{j}^{*}X_{j}X_{1}-X_{j}^{*}X_{j}d_{1})v,v)+((X_{0}-X_{0}^{*})v,v)$ (2.33)
@@
$-\sum_{j=1}^{N}\left(\Big{(}(d_{1}X_{j}^{*}X_{j}+2(X_{j}d_{1})X_{j}+d_{1}d_{j}X_{j}+(X_{j}^{*}X_{j}d_{1})+(X_{j}d_{1})d_{j})\Big{)}v,v\right)-(d_{0}v,v)$ (2.35)
$=-\sum_{j,k=1}^{N}(2c^{j1}_{k}X_{k}X_{j}v,v)-d_{1}\sum_{j=1}^{N}X_{j}^{*}X_{j}-\sum_{j,k,l=1}^{N}(c^{jk}_{l}X_{l}v,v)-\sum_{j,k=1}^{N}((X_{j}c_{k}^{j1})X_{k}v,v)$ (2.36)
$-\sum_{j,k=1}^{N}(d_{j}c_{k}^{j1}X_{k}v,v)-\sum_{j=1}^{n}\left(\Big{(}-(X_{1}d_{j})+2(X_{j}d_{1})+d_{1}d_{j}\Big{)}X_{j}v,v\right)$ (2.37)
$-\sum_{j=1}^{N}\left(((X_{j}^{*}X_{j}d_{1})+(X_{j}d_{1})d_{j})\,v,v\right)-(d_{0}v,v)$ (2.38)
. (2.39)
@@
the term containing $L_{2}$, recalling that $f$ is a real function and that $X_{1}f$ takes purely imaginary values, we obtain
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)\geq\sum_{j=1}^{N}(Q_{j}e^{\lambda f}u,e^{\lambda f}u)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}v,v)-\lambda^{2}C_{K}\|v\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K),$ (2.53)
@@
$\mathrm{Im}(e^{\lambda f}P_{1}^{*}u,e^{\lambda f}u)$ (2.56)
$\geq\small{\sum_{j=1}^{N}}(Q^{\prime}_{j}e^{\lambda f}u,e^{\lambda f}u)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}e^{\lambda f}u,e^{\lambda f}u)-\lambda^{2}C_{K}\|e^{\lambda f}u\|_{L^{2}}^{2}$ (2.57)
@@
$\mathrm{Im}(e^{\lambda f}P_{1}^{*}u,e^{\lambda f}u)$ $=\mathrm{Im}({P_{1}^{*}}^{f}e^{\lambda f}u,e^{\lambda f}u),$ (2.64)
@@
${P_{1}^{*}}^{f}=$ $e^{\lambda f}P_{1}^{*}e^{-\lambda f}=P_{1}^{*}(x,D+\lambda Df)$ (2.65)
@@
    $\|e^{\lambda f}P_{1}^{*}u\|_{L^{2}}^{2},\|e^{\lambda f}P_{1}u\|_{L^{2}}^{2}\geq C\lambda\|e^{\lambda f}u\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K).$ (2.87)
@@
    $\|e^{\lambda f}P_{1}^{*}u\|_{L^{2}}^{2},\|e^{\lambda f}P_{1}u\|_{L^{2}}^{2}\geq C\lambda\|e^{\lambda f}u\|_{H^{\frac{1}{r}}}^{2},\quad\forall u\in C_{0}^{\infty}(K).$ (2.88)
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)$ (2.89)
$\geq\small{\sum_{j=1}^{N}}(Q_{j}e^{\lambda f}u,e^{\lambda f}u)+\lambda^{3}\|\,|X_{1}f|^{\frac{3}{2}}v\|_{L^{2}}^{2}-\lambda^{2}C_{K}\|v\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K),$ (2.90)
@@
where $ic^{j1}_{k},ic^{\prime}_{k}$, $k=1,\ldots,N$, are smooth and real valued functions. Therefore, once $\lambda$ is big enough, the proof of $(i)$ and of $(ii)$ depends on the estimate we can prove for the second order term $\small{\sum_{j=1}^{n}}\mathrm{Re}(Q_{j}e^{\lambda f}u,e^{\lambda f}u)$. Then, let us focus on this term and show how different hypotheses lead to different lower bounds.
In the sequel we will use again the notation $v:=e^{\lambda f}u$. First observe that, by Cauchy-Schwartz inequality,
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{3}\|\,|X_{1}f|^{\frac{3}{2}}v\|_{L^{2}}^{2}-\lambda^{2}C_{K}\|v\|_{L^{2}}^{2},$ (2.100)
@@
$\mathrm{Im}(e^{\lambda f}P_{1}u,e^{\lambda f}u)\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{2}\frac{C_{K}}{2}\|v\|_{L^{2}}^{2},$ (2.101)
@@
Finally, by Cauchy-Schwartz inequality,
$\|e^{\lambda f}P_{1}u\|_{L^{2}}^{2}\geq 2\delta\Big{(}\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+(\lambda^{2}\frac{C_{K}}{2}-\frac{\delta}{2})\|v\|_{L^{2}}^{2}\Big{)},$ (2.102)
which gives, taking $\delta$ small enough and not necessarily depending on $K$ (since we can always assume $\frac{C_{K}}{2}\cdot\lambda>1$ and take, for instance $\delta=1/2<\lambda^{2}\frac{C_{K}}{4}$), we reach
$\|e^{\lambda f}P_{1}u\|_{L^{2}}^{2}\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{2}\frac{C_{K}}{4}\|v\|_{L^{2}}^{2},$ (2.103)
@@
The solvability estimate, Hahn-Banach and Riesz representation theorems, alltogether give the result (see, for instance, Lemma 1.2.30 in [35]). Since $X_{1}$ is nondegenerate (see also Remark 2.1), given $C_{0}>0$ and $x_{0}\in\mathbb{R}^{n+1}$, it is always possible to find a smooth function $f$ and a compact set $K\subset\mathbb{R}^{n+1}$ containing $x_{0}$ in its interior, such that $iX_{1}f(x)\geq C_{0}>0$ for all $x\in K$. Hence, by the Carleman estimate for $P_{1}^{*}$, for $x_{0}\in\mathbb{R}^{n+1}$ and for the compact $K$ chosen as above and containing $x_{0}$ in its interior, there exists $\lambda_{0}(K)>1$ and $C>0$ such that, for all $\lambda>\lambda_{0}$,
$\|e^{\lambda f}P_{1}^{*}\|_{L^{2}}\geq C\|e^{\lambda f}u\|_{H^{-s}},\quad\forall u\in C_{0}^{\infty}(K),$ (2.104)
where $s=0$ in case (i) and $s=-1/r$ in case (ii). By estimating from above and from below $e^{\lambda f}$ on $K$ with positive constants, (2.104) easily implies the solvability estimate with $s=0$ under the hypothesis in part (i), and with $s=-1/r$ under the hypothesis in part (ii). From the solvability estimate, by the standard considerations mentioned at the beginning of the proof (see [35]) we get the $L^{2}-L^{2}$ and $H^{-\frac{1}{r}}-L^{2}$ local solvability of $P_{1}$ at $x_{0}$ according to the hypotheses in (i) and (ii) respectively. Finally, by the arbitrariness of $x_{0}$, we conclude the local solvability of $P_{1}$ at any point of $\mathbb{R}^{n+1}$.
@@
When the vector $X_{1}$ field is nondegenerate only on a compact $K\subset\mathbb{R}^{n+1}$, using the previous strategy we can reach the local solvability of $P_{1}$ and of $P_{1}^{*}$ at any point of the interior $U_{K}$ of $K$.

### 08 Apr 2024 v2

@@ 2. Carleman estimate and local solvability for $P_{1}$ and $P_{1}^{*}$
    $[X_{j},X_{k}](x,D)=\sum_{l=1}^{N}c^{jk}_{l}(x)X_{l}(x,D),\quad\forall x\in\mathbb{R}^{n+1},$ (2.4)
    where $[X,Y]:=XY-YX$ denotes the commutator of the operators $X$ and $Y$.
@@
Notations. Below $N$ will always be a positive integer, $\lambda$ will be a real number, whereas $A=\{a_{ji}(x)\}_{\begin{subarray}{c}j=0,\ldots,N\\ i=1,\dots,n+1\end{subarray}}$ will denote the variable coefficient matrix containing the coefficients of the vector fields $iX_{j}$, for $j=0,\ldots,N$. For a matrix $A$ as above and $\alpha\in\mathbb{N}^{n+1}\bigcup\{0\}^{n+1}$, we define
$\|\nabla^{\alpha}A\|_{L^{\infty}(\Omega)}:=\sup_{\begin{subarray}{c}j=1,\ldots,N\\ i=1,\dots,n+1\end{subarray}}\|\partial^{\alpha}a_{ji}\|_{L^{\infty}(\Omega)},\quad\forall\Omega\subset\mathbb{R}^{n+1},$ (2.8)
where $\partial^{\alpha}f=\partial_{1}^{\alpha_{1}}\ldots\partial_{n+1}^{\alpha_{n+1}}$ Throughout the paper we will write $\|\cdot\|_{L^{\infty}}$, $\|\cdot\|_{L^{2}}$, $\|\cdot\|_{H^{s}}$ without specifying the set $\Omega$ where the $L^{\infty}(\Omega)$, $L^{2}(\Omega)$, and the $H^{s}(\Omega)$ norms are taken. Given that we will be working with compactly supported functions, the set $\Omega$ will be the fixed support of the function (or its interior), so we omit it for simplicity. However, note that we can take $\|\nabla^{\alpha}A\|_{L^{\infty}}=\|\nabla^{\alpha}A\|_{L^{\infty}(\mathbb{R}^{n+1})}$ in the coming estimates when $A$ has $C^{\infty}_{b}(\mathbb{R}^{n+1})$ coefficients. Finally, we shall use the notation $\Gamma(\{X_{j}\}_{j=1}^{N})$ for the distribution spanned by the system of vector fields $\{X_{j}\}_{j=1}^{N}$.
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)$ (2.9)
$\geq\small{\sum_{j=1}^{N}}(Q_{j}e^{-\lambda f}u,e^{-\lambda f}u)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}e^{-\lambda f}u,e^{-\lambda f}u)-\lambda^{2}C_{K}\|e^{-\lambda f}u\|_{L^{2}}^{2},$ (2.10)
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u).$ (2.13)
@@
$e^{-\lambda f}P_{1}(x,D)u=e^{-\lambda f}P_{1}(x,D)e^{\lambda f}e^{-\lambda f}u=P^{f}_{1}(x,D)e^{-\lambda f}u,$ (2.14)
@@
where $L_{2},L_{1},L_{0}$ are the partial differential differential operators of order 2, 1, and 0, given, respectively, by
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)$ $=\mathrm{Im}(P_{1}e^{-\lambda f}u,e^{-\lambda f}u)+\mathrm{Im}(L_{2}e^{-\lambda f}u,e^{-\lambda f}u)+\mathrm{Im}(L_{1}e^{-\lambda f}u,e^{-\lambda f}u)+\mathrm{Im}(L_{0}e^{-\lambda f}u,e^{-\lambda f}u),$ (2.29)
so, for simplicity, we handle each term on the right-hand side separately.
@@
Now, on putting $v:=e^{-\lambda f}u$, and using the previous observation, we have
$2i\mathrm{Im}(P_{1}e^{-\lambda f}u,e^{-\lambda f}u)$ $=2i\sum_{j=1}^{N}\mathrm{Im}(X_{1}X_{j}^{*}X_{j}v,v)+2i\,\mathrm{Im}(X_{0}v,v)$ (2.32)
$=\sum_{j=1}^{N}\left((X_{1}X_{j}^{*}X_{j}-X_{j}^{*}X_{j}X_{1}-X_{j}^{*}X_{j}d_{1})v,v\right)+\left((X_{0}-X_{0}^{*})v,v\right)$ (2.33)
@@
$-\sum_{j=1}^{N}\Bigg{(}\Big{(}d_{1}X_{j}^{*}X_{j}+2(X_{j}d_{1})X_{j}+d_{1}d_{j}X_{j}+(X_{j}^{*}X_{j}d_{1})+(X_{j}d_{1})d_{j}\Big{)}v,v\Bigg{)}-(d_{0}v,v)$ (2.35)
$=-\sum_{j,k=1}^{N}(2c^{j1}_{k}X_{k}X_{j}v,v)-d_{1}\sum_{j=1}^{N}X_{j}^{*}X_{j}-\sum_{j,k,l=1}^{N}(c^{jk}_{l}X_{l}v,v)-\sum_{j,k=1}^{N}\left((X_{j}c_{k}^{j1})X_{k}v,v\right)$ (2.36)
$-\sum_{j,k=1}^{N}(d_{j}c_{k}^{j1}X_{k}v,v)-\sum_{j=1}^{n}\Bigg{(}\Big{(}-(X_{1}d_{j})+2(X_{j}d_{1})+d_{1}d_{j}\Big{)}X_{j}v,v\Bigg{)}$ (2.37)
$-\sum_{j=1}^{N}\Bigg{(}\big{(}(X_{j}^{*}X_{j}d_{1})+(X_{j}d_{1})d_{j}\big{)}\,v,v\Bigg{)}-(d_{0}v,v).$ (2.38)
@@
As for the term containing $L_{2}$, recalling that $f$ is a real function and that $X_{1}f$ takes purely imaginary values, we obtain
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)\geq\sum_{j=1}^{N}(Q_{j}v,v)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}v,v)-\lambda^{2}C_{K}\|v\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K),$ (2.53)
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}^{*}u,e^{-\lambda f}u)$ (2.56)
$\geq\small{\sum_{j=1}^{N}}(Q^{\prime}_{j}e^{-\lambda f}u,e^{-\lambda f}u)+\lambda^{3}\sum_{j=1}^{N}((-iX_{1}f)|X_{j}f|^{2}e^{-\lambda f}u,e^{-\lambda f}u)-\lambda^{2}C_{K}\|e^{-\lambda f}u\|_{L^{2}}^{2}$ (2.57)
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}^{*}u,e^{-\lambda f}u)$ $=\mathrm{Im}({P_{1}^{*}}^{f}e^{-\lambda f}u,e^{-\lambda f}u),$ (2.64)
@@
${P_{1}^{*}}^{f}=$ $e^{-\lambda f}P_{1}^{*}e^{\lambda f}=P_{1}^{*}(x,D+\lambda Df)$ (2.65)
@@
    $\|e^{-\lambda f}P_{1}^{*}u\|_{L^{2}}^{2},\|e^{-\lambda f}P_{1}u\|_{L^{2}}^{2}\geq C\lambda\|e^{-\lambda f}u\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K).$ (2.87)
@@
    $\|e^{-\lambda f}P_{1}^{*}u\|_{L^{2}}^{2},\|e^{-\lambda f}P_{1}u\|_{L^{2}}^{2}\geq C\lambda\|e^{-\lambda f}u\|_{H^{\frac{1}{r}}}^{2},\quad\forall u\in C_{0}^{\infty}(K).$ (2.88)
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)$ (2.89)
$\geq\small{\sum_{j=1}^{N}}(Q_{j}e^{-\lambda f}u,e^{-\lambda f}u)+\lambda^{3}\|\,|X_{1}f|^{\frac{3}{2}}e^{-\lambda f}u\|_{L^{2}}^{2}-\lambda^{2}C_{K}\|e^{-\lambda f}u\|_{L^{2}}^{2},\quad\forall u\in C_{0}^{\infty}(K),$ (2.90)
@@
where $ic^{j1}_{k},ic^{\prime}_{k}$, $k=1,\ldots,N$, are smooth and real valued functions. Therefore, once $\lambda$ is big enough, the proof of $(i)$ and of $(ii)$ depends on the estimate we can prove for the second order term $\small{\sum_{j=1}^{n}}\mathrm{Re}(Q_{j}e^{-\lambda f}u,e^{-\lambda f}u)$. Then, let us focus on this term and show how different hypotheses lead to different lower bounds.
In the sequel we will use again the notation $v:=e^{-\lambda f}u$. First observe that, by Cauchy-Schwarz inequality,
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{3}\|\,|X_{1}f|^{\frac{3}{2}}v\|_{L^{2}}^{2}-\lambda^{2}C_{K}\|v\|_{L^{2}}^{2},$ (2.100)
@@
$\mathrm{Im}(e^{-\lambda f}P_{1}u,e^{-\lambda f}u)\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{2}\frac{C_{K}}{2}\|v\|_{L^{2}}^{2},$ (2.101)
@@
Finally, by Cauchy-Schwarz inequality,
$\|e^{-\lambda f}P_{1}u\|_{L^{2}}^{2}\geq 2\delta\Big{(}\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+(\lambda^{2}\frac{C_{K}}{2}-\frac{\delta}{2})\|v\|_{L^{2}}^{2}\Big{)},$ (2.102)
which gives, taking $\delta$ small enough and not necessarily depending on $K$ (since we can always assume $\frac{C_{K}}{2}\cdot\lambda>1$ and take, for instance, $\delta=1/2<\lambda^{2}\frac{C_{K}}{4}$), we reach
$\|e^{-\lambda f}P_{1}u\|_{L^{2}}^{2}\geq\lambda\frac{C_{0}}{2}\small{\sum_{j=1}^{N}}\|X_{j}v\|_{L^{2}}^{2}+\lambda^{2}\frac{C_{K}}{4}\|v\|_{L^{2}}^{2},$ (2.103)
@@
The solvability estimate, Hahn-Banach and Riesz representation theorems, alltogether give the result (see, for instance, Lemma 1.2.30 in [35]). Since $X_{1}$ is nondegenerate (see also Remark 2.1), given $C_{0}>0$ and $x_{0}\in\mathbb{R}^{n+1}$, it is always possible to find a smooth function $f$ and a compact set $K\subset\mathbb{R}^{n+1}$ containing $x_{0}$ in its interior, such that $iX_{1}f(x)\geq C_{0}>0$ for all $x\in K$. Hence, by the Carleman estimate for $P_{1}^{*}$, for $x_{0}\in\mathbb{R}^{n+1}$, and for the compact $K$ chosen as above and containing $x_{0}$ in its interior, there exists $\lambda_{0}(K)>1$ and $C>0$ such that, for all $\lambda>\lambda_{0}$,
$\|e^{-\lambda f}P_{1}^{*}\|_{L^{2}}\geq C\|e^{-\lambda f}u\|_{H^{-s}},\quad\forall u\in C_{0}^{\infty}(K),$ (2.104)
where $s=0$ in case (i) and $s=-1/r$ in case (ii). By estimating from above and from below $e^{-\lambda f}$ on $K$ with positive constants, (2.104) easily implies the solvability estimate with $s=0$ under the hypothesis in part (i), and with $s=-1/r$ under the hypothesis in part (ii). From the solvability estimate, by the standard considerations mentioned at the beginning of the proof (see [35]) we get the $L^{2}-L^{2}$ and $H^{-\frac{1}{r}}-L^{2}$ local solvability of $P_{1}$ at $x_{0}$ according to the hypotheses in (i) and (ii) respectively. Finally, by the arbitrariness of $x_{0}$, we conclude the local solvability of $P_{1}$ at any point of $\mathbb{R}^{n+1}$.
@@
When the vector field $iX_{1}$ is nondegenerate only on a compact $K\subset\mathbb{R}^{n+1}$, using the previous strategy we can reach the local solvability of $P_{1}$ and of $P_{1}^{*}$ at any point of the interior $U_{K}$ of $K$.